### v1

# GIT-BO: High-Dimensional Bayesian Optimization with Tabular Foundation Models

Abstract
Bayesian optimization (BO) effectively optimizes expensive black-box functions but faces significant challenges in high-dimensional spaces (dimensions exceeding 100) due to the curse of dimensionality. Existing high-dimensional BO methods typically leverage low-dimensional embeddings or structural assumptions to mitigate this challenge, yet these approaches frequently incur considerable computational overhead and rigidity due to iterative surrogate retraining and fixed assumptions. To address these limitations, we propose Gradient-Informed Bayesian Optimization using Tabular Foundation Models (GIT-BO), an approach that utilizes a pre-trained tabular foundation model (TFM) as a surrogate, leveraging its gradient information to adaptively identify low-dimensional subspaces for optimization. We propose a way to exploit internal gradient computations from the TFM’s forward pass by creating a gradient-informed diagnostic matrix that reveals the most sensitive directions of the TFM’s predictions, enabling optimization in a continuously re-estimated active subspace without the need for repeated model retraining. Extensive empirical evaluation across 23 synthetic and real-world benchmarks demonstrates that GIT-BO consistently outperforms four state-of-the-art Gaussian process-based high-dimensional BO methods, showing superior scalability and optimization performances especially as dimensionality increases up to 500 dimensions. This work establishes foundation models, augmented with gradient-informed adaptive subspace identification, as highly competitive alternatives to traditional Gaussian process-based approaches for high-dimensional Bayesian optimization tasks.
1 Introduction
Bayesian optimization (BO) is widely recognized as an effective and sample-efficient technique for black-box optimization, essential in machine learning (dewancker2016bayesian, ; snoek2012practical, ), engineering design (kumar2024reverse, ; wang2022bayesian, ; zhang2020bayesian, ; yu2025fast, ), and hyperparameter tuning (klein2017fast, ; wu2019hyperparameter, ). However, BO can struggle in high-dimensional spaces (especially when dimension $D>$ 100) due to the curse of dimensionality, making it difficult for conventional Gaussian process (GP)-based methods to effectively explore complex objective functions (malu2021bayesian, ; wang2023recent, ; hvarfner2024vanilla, ; xu2024standard, ; letham2020re, ). Existing high-dimensional BO approaches mitigate these challenges by exploiting low-dimensional subspaces, but still rely heavily on iterative model retraining and structural assumptions, incurring substantial computational overhead (rana2017high, ; eriksson2019global, ; letham2020re, ; eriksson2021high, ).
Recent advances in tabular foundation models (TFM) have proposed using Prior-Data Fitted Networks (PFNs) to bypass surrogate refitting in Bayesian Optimization. Initially demonstrated for simple low-dimensional BO problems (muller2023pfns, ), the PFN-based BO was further validated for constrained engineering problems yu2025fast , using the parallel processing capabilities of the transformer architecture to outperform other methods in both speed and performance. TabPFN v2, a recent TFM model, extends the capabilities of TabPFN to handle inputs up to 500 dimensions, demonstrating superior performance in zero-shot classification, regression, and time series prediction tasks (hollmann2025accurate, ).
While TabPFN v2 presents an exciting opportunity as a high-dimensional BO surrogate due to its computational efficiency and predictive accuracy, its fixed-parameter foundation model architecture restricts its ability to dynamically adapt kernel hyperparameters. This is in contrast to traditional GP models, which adjust their kernel parameters during optimization to effectively identify principal searching directions in high-dimensional spaces (eriksson2021high, ; hvarfner2024vanilla, ; xu2024standard, ). Addressing this limitation requires an approach that can leverage the strengths of foundation models while introducing adaptive mechanisms for subspace exploration, a gap we aim to bridge in this work.
To address this, we introduce Gradient-Informed Tabular Bayesian Optimization (GIT-BO), a framework that utilizes gradient information from the foundation model’s forward pass to dynamically guide adaptive subspace exploration. Specifically, our contributions are as follows:
  • We propose GIT-BO, a novel Bayesian optimization approach that employs a tabular foundation model as a surrogate, demonstrating superior scaling properties and performance on complex optimization landscapes where traditional GP-based methods struggle.
  • Our gradient-informed subspace identification technique adaptively learns promising search directions in high-dimensional spaces, exploiting the foundation model’s adaptive gradient knowledge during the optimization trajectory.
  • We compare GIT-BO against four popular state-of-the-art (SOTA) GP high-dimensional BO methods across twenty-three diverse benchmarks, including synthetic functions and real-world problems, introducing foundation model surrogates as a viable alternative for complex Bayesian optimization tasks.
The rest of the paper is structured as follows. Section 2 states the background problem and discusses related work. Section 3 presents the GIT-BO algorithm. Section 4 defines the benchmark high-dimensional BO algorithms of interest, the test problem set, and the evaluation methods for algorithm performance. The results are presented in Section 5, and Section 6 is the concluding discussion.
2 Background
Bayesian Optimization and the High-Dimensional Challenge
Bayesian optimization is a sample-efficient approach for optimizing expensive black-box functions where the objective is to find $x^{*}\in\arg\min_{x\in\mathcal{X}}f(x)$ with $\mathcal{X}=[0,1]^{D}$. The algorithm leverages a surrogate model, typically a GP, to approximate the objective function, and calculate the acquisition function to determine the next query points until it converges to the optimum. As dimensionality $D$ grows, effectively exploring the space often requires more observations $n$, pushing GPs into regimes of high computational cost ($O(n^{3})$ complexity) (liu2020gaussian, ; wang2023recent, ; santoni2024comparison, ).
Existing Strategies for Tackling High-Dimensional Bayesian Optimization Challenge
Several baseline approaches have been developed to address high-dimensional BO challenges. Random Embedding Bayesian Optimization (REMBO) projects the optimization domain onto a randomly selected low-dimensional linear subspace (wang2016bayesian, ; nayebi2019framework, ; letham2020re, ; papenmeier2022increasing, ; ziomek2023random, ). However, REMBO suffers from distorted objective values due to clipping points back into the feasible domain, impairing GP calibration (wang2016bayesian, ). Hashing-Enhanced Subspace Bayesian Optimization (HESBO) replaces Gaussian projections with a sparse hashing sketch that keeps every point feasible, at the price of a slightly smaller chance of containing the active subspace (nayebi2019framework, ). Another approach, Adaptive Linear Embedding Bayesian Optimization (ALEBO), creates a more effective optimization framework by employing linear constraints that preserve the embedding’s linearity to further improve optimization performance (letham2020re, ). All three methods, however, require the embedding dimension $d$ to be defined by users. BAXUS lifts this restriction by growing a nested sequence of random sub-spaces as data accumulate (papenmeier2022increasing, ).
A complementary line of work models the full space but exploits structure. TURBO maintains multiple GP-driven trust regions that expand when progress is made and shrink otherwise, achieving strong results up to 200 variables with thousands of evaluations (eriksson2019global, ). SAASBO places half-Cauchy shrinkage priors on GP length-scales, automatically selecting a sparse axis-aligned subset of salient variables (eriksson2021high, ). However, its computational cost grows cubically with the sample count rather than the dimension. While these methods offer significant improvements, they still entail limitations such as fixed-dimensional embedding choices or incremental computational overhead, highlighting the need for alternative approaches that can more dynamically and efficiently address high-dimensional optimization problems.
Foundation Models as Bayesian Optimization Surrogates
Tabular foundation models have emerged as promising alternatives to GP-based methods in Bayesian optimization. Prior-data Fitted Networks (PFNs) (muller2021transformers, ) demonstrated that pre-trained tabular foundation models can effectively replace GPs as surrogate models. PFNs4BO (muller2023pfns, ) showed competitive performance on low-dimensional BO tasks with significantly reduced computational overhead compared to GPs. However, PFNs inherently lack the adaptive capability of GPs in dynamically updating kernel hyperparameters during optimization, limiting their ability to discover low-rank structures within high-dimensional spaces.
Interestingly, a single forward pass yields internal gradients with respect to inputs and hidden states. Studies in the field of Large Language Models (LLMs) show that leveraging gradient information can help guide model behavior (cai2024self, ) or perform rapid adaptation to distribution shifts (deng2025learning, ). These findings motivate our exploration of gradient-informed techniques within Bayesian optimization, particularly to harness the intrinsic structure of high-dimensional search spaces.
Gradient-Enhanced Bayesian Optimization
Previous work demonstrates how derivative information can accelerate Bayesian optimization through various mechanisms or capture the search space structure (ahmed2016we, ; wu2016parallel, ; wei2025gradient, ). When explicit derivatives are available, a direct integration of gradient information is to jointly model function values and gradients in a Gaussian process (ahmed2016we, ). Strategies such as the parallel knowledge gradient method for batch BO (wu2016parallel, ) and sampling the most informative observation subset using GP log-likelihood gradients (wei2025gradient, ) have shown faster convergence of BO. However, these methods inherently depend on precise GP modeling, which becomes less effective as dimensionality grows.
To overcome these limitations, we introduced the Gradient-Informed Bayesian Optimization using Tabular Foundation Models (GIT-BO) algorithm—a method that does not require fitting a GP. Our approach fundamentally departs from the GP gradient-enhanced BO methods: We combine the gradient information obtained directly from the forward pass of tabular foundation models and integrate subspace identification, an established strategy in high-dimensional optimization. We can approximate high-dimensional target measures with low-dimensional updates by adapting the principal feature detection framework based on $\phi$-Sobolev inequalities (li2024principal, ). This approach combines the computational efficiency of foundation models with subspace identification capabilities, enabling effective exploration in high-dimensional spaces.
3 The GIT-BO Algorithm
We now present our GIT-BO high-dimensional BO framework. The framework consists of four main components: the surrogate model (TabPFN v2), the gradient-based subspace identification, the choice of acquisition function, and the algorithm that ties these together.
Surrogate Modeling with TabPFN
We use TabPFN, specifically the 500-dimensional TabPFN v2 TFM model from (hollmann2025accurate, ), as our surrogate model to predict the objective function values using in-context learning. The optimization objective is defined as minimizing an unknown function $f(x)$ over a domain $x\in\mathcal{X}\subset\mathbb{R}^{D}$. At any BO iteration, let $D_{n}={(x_{i},y_{i})}_{i=1}^{n}$ be the dataset of points sampled so far and their observed function values $y_{i}=f(x_{i})$. In a standard BO setting, one would fit a GP posterior $p(f|\mathcal{D}n)$ to these data. Here instead, we leverage TabPFN to obtain a predictive model. We provide the dataset $D_{n}$ to TabPFN as input context (taking a set of labeled examples as part of its input sequence) and query the model for its prediction at any candidate point $x$. TabPFN then returns an approximate posterior predictive distribution $\text{PFN}(y\mid x,D_{n})$ informed by both the new point $x$ and the context of observed samples. In practice, TabPFN produces a prediction in a single forward pass as an approximation to the Bayesian posterior mean $\mu_{n}(x)$ for $f(x)$ as this predictive mean given the provided data. In our method, we review the previous series of PFN work (pmlr-v202-muller23a, ) and implement the calculation of predictive mean $\mu_{n}(x)$ and predictive variance $\sigma_{n}^{2}(x)$ with the latest TabPFN v2 regressor model.
Gradient-Informed Subspace Identification and Sampling
For active subspace identification, we leverage gradient information $\nabla_{x}\mu_{n}(x)$ obtained through one-step backpropagation on TabPFN’s predictive mean values, which naturally varies across data points at each iteration. Inspired by studies on dimension reduction in nonlinear Bayesian inverse problems, which employs techniques $\phi$-Sobolev inequalities and gradient-based approaches, we approximate the diagnostic (Fisher information) matrix ${H}=\mathbb{E}_{\mu}[\nabla_{x}\mu_{n}(x)\,\nabla_{x}\mu_{n}(x)^{\top}]$ zahm2022certified ; li2024sharp ; li2024principal ; ly2017tutorial . In this study of high-dimensional problems with only $D>100$, we select the top $r\;(D>>r)$ dominant eigenvectors in ${H}$ as the principal vectors ${V}_{r}$ that span our gradient-informed active subspace (GI subspace). Next, we restrict the next exploration to this subspace and map the search candidates back to the original space by $x_{\text{cand}}=\bar{x}_{ref}+{V}_{r}{z},\;{z}\sim\mathrm{U}([-1,1]^{r})$, where the reference point $\bar{x}_{ref}$ is the centroid of the observed data. Finally, we pass $x_{\text{cand}}$ to the acquisition function for determining the next sample to evaluate. In GIT-BO, we recalculate the GI subspace at each iteration as the subspace changes when observed data increases. For computational practicality and exploration-exploitation balance, we default to $r=15$ in GIT-BO; a detailed sensitivity analysis on the choice of $r$ is presented in the Appendix.
Acquisition Function
We implement the Thompson Sampling (TS) acquisition as our acquisition function due to its previous success in high-dimensional BO (eriksson2019global, ). For TS, we approximate it by sampling from the predictive distribution at each sample $x$ and draw a fixed number of 512 random samples $\tilde{f}(\cdot)$ from the surrogate’s posterior. The candidate $x_{\text{next}}$ with the highest sampled value is selected as the next query $\tilde{f}(x_{\text{next}})$. This leverages the full predictive uncertainty and tends to balance exploration and exploitation implicitly (shahriari2015taking, ; eriksson2019global, ). The key difference here is that we restrict $x$ to our learned GI subspace. After selecting the next query point $x_{\text{next}}$, we evaluate the true objective function to obtain $y_{\text{next}}$.
Putting everything together, Figure 1 and Algorithm 1 outlines the GIT-BO procedure combining TabPFN with gradient-informed subspace search. The implementation details of GIT-BO are stated in the Appendix.
Figure 1: GIT-BO algorithm overview. The method operates in four stages: (1) Initial observed samples are collected in the high-dimensional space $\mathbb{R}^{D}$; (2) TabPFN v2, a fixed-weight tabular foundation model generates predictions of the objective space at inference time using in-context learning; (3) Gradient information from TabPFN’s forward pass ($\nabla\hat{\mu}(x)$) is used to identify a low-dimensional gradient-informed (GI) subspace. The predicted mean and variance are used for acquisition value calculations $\hat{\mu}(x),\hat{\sigma}^{2}(x)$; (4) The next sample point ($x_{\text{next}}$) is selected within this GI subspace with the highest acquisition value and appended to the dataset for iterative search until stopping criteria is met.
Algorithm 1 Gradient-Informed Bayesian Optimization using Tabular Foundation Models (GIT-BO)
0:  objective $f$, domain $\mathcal{X}\!\subset\!\mathbb{R}^{D}$, init. size $n_{0}$, iteration budget $I$, subspace dim. $r$
1:  Draw $n_{0}$ LHS points $x_{i}$ and set $y_{i}=f(x_{i})$;  $D\leftarrow\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$
2:  for $i=1\;\text{to}\;I$ do
3:     Fit TabPFN on $D$ to obtain mean $\mu_{n}$ and variance $\sigma^{2}_{n}$
4:     Calculate gradient $\nabla_{x}\mu_{n}(x_{i})$ for all $(x_{i},y_{i})\!\in\!D$
5:     Form diagnostic matrix ${H}=\mathbb{E}_{\mu}[\nabla_{x}\mu_{n}(x)\,\nabla_{x}\mu_{n}(x)^{\top}]$;
6:     ${V}_{r}\leftarrow$ top-$r$ eigvectors of ${H}$
7:     Set $x_{\text{cand}}\leftarrow x_{\text{ref}}+V_{r}z$, with $x_{\text{ref}}\leftarrow\bar{x}$ and $z\sim\mathrm{U}([-1,1]^{r})$
8:     $x_{\text{next}}\leftarrow\arg\max_{j}\mathrm{ThompsonSampling}_{n}(x_{\text{cand}})$
9:     Evaluate $y_{\text{next}}=f(x_{\text{next}})$and append the query point data $D\leftarrow D\cup\{(x_{\text{next}},y_{\text{next}})\}$
10:  end for
11:  return $x^{\star}=\arg\min_{(x,y)\in D}y$
4 Experiment
This section evaluates GIT-BO’s performance on a diverse collection of synthetic optimization benchmarks and real-world engineering problems.
4.1 Experiment Setups
This section outlines our empirical approach to evaluating and comparing different high-dimensional Bayesian optimization algorithms, specifically focusing on assessing their performance across various complex synthetic and engineering benchmarks.
Benchmark Algorithms
We benchmark GIT-BO against random search (bergstra2012random, ) and four high-dimensional BO methods, including SAASBO (eriksson2021high, ), TURBO (eriksson2019global, ), HESBO (nayebi2019framework, ), and ALEBO (letham2020re, ). These high-dimensional GP BO algorithms have been introduced in Section 2. For HESBO and ALEBO, we varied $d_{\text{embedding}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO (d={10, 20}). The implementation of SAASBO and TURBO is taken from BoTorch tutorials (balandat2020botorch, ). The implementation of HESBO and ALEBO are taken from their original published paper and code (nayebi2019framework, ; letham2020re, ). Additional details of the algorithm implementation are listed in the Appendix.
Test Problems
This study incorporates a diverse set of high-dimensional optimization problems, including 11 synthetic problems and 12 real-world benchmarks. Synthetic and scalable problems include: Ackley, Rosenbrock, Dixon-Price, Levy, Powell, Griewank, Rastrigin, Styblin-Tang, and Michalewicz. Additionally, we tested on two Synthetic problems in LassoBench (LassoSyntheticMedium and LassoSyntheticHigh) and one real-world hyperparameter optimization (HPO) problem LassoDNA. The rest of the application problems are collected from previous optimization studies and conference benchmarks: the power system optimization problems in CEC2020, Rover, SVM HPO, MOPTA08 car problem, and two Mazda car problems. As this study focus on the high-dimensional characteristic of problem, we make all our benchmark problems single and unconstrained for testing. Therefore, we have applied penalty transforms to all real-world problems with constraints and perform average weighting to the two multi-objective Mazda problems. The details of benchmark transformation and implementation are detailed in the Appendix. Among the 23 benchmarks, 10 (Synthetic + Rover) are scalable problems. To evaluate the algorithms’ performance with respect to dimensionality, we solve the scalable problems for $D=\{100,200,300,400,500\}$. Therefore, we have experimented with a total of $5\times 10+13=63$ different variants of the benchmark problems. The benchmarks’ details and implementations are listed in the Appendix.
Algorithm Tests
The algorithm evaluation aims to thoroughly compare GIT-BO to current SOTA Bayesian optimization techniques. This study focuses on minimizing the objective function for the given test problems. For each test problem, our experiment consists of 20 independent trials, each utilizing a distinct random seed. To ensure fair comparison, we initialize each algorithm with an identical set of 200 samples, generated through Latin Hypercube Sampling with consistent random seeds across all trials. During each iteration, each algorithm selects one sample to evaluate next.
To execute this extensive benchmarking process, we utilized a distributed server infrastructure featuring Intel Xeon Platinum 8480+ CPUs and NVIDIA H100 GPUs. Individual experiments were conducted with the same amount of compute allocated: a single H100 GPU node with 24 CPU cores and 250GB RAM.
4.2 Evaluation Metrics
Optimization Fixed-budget Convergence Analysis
Fixed-budget evaluations is a technique for comparing the efficiency of optimization algorithms by allotting specific computational resources for their execution (hansen2022anytime, ). We apply a “fixed-iteration” budget approach in our study, initially running all algorithms (GIT-BO, TurBO, HESBO, SAASBO, and ALEBO) for 100 iterations. Additionally, we run GIT-BO, TurBO, and HESBO for an extra 100 iterations (total of 200 iterations) because these methods are roughly two orders of magnitude faster than SAASBO and ALEBO, allowing us to explore their long-term convergence behavior more thoroughly.
Statistical Ranking
For comprehensively compare and evaluate the performance of the Bayesian optimization algorithms, statistical ranking techniques are employed instead of direct performance measurements of the optimization outcome. In this study, we define the optimization performance result as the median of the minimum result found (final incumbent) across the 20 optimization trials of each algorithm. By statistically ranking the results, we were able to standardize the comparisons across different problems, since various optimization challenges can produce objective values of vastly different magnitudes. Furthermore, using this ranking allowed us to reduce the distorting effects of unusual or extreme data points that might influence our evaluation.
We conduct our statistical analysis using the Friedman and Wilcoxon signed-rank tests, complemented by Holm’s alpha correction. These non-parametric approaches excel at processing benchmarking result data without assuming specific distributions, which is critical for handling optimization results with outliers. These statistical methods effectively handle the dependencies in our setup, where we used the same initial samples and seeds to test all algorithms. The Wilcoxon signed-rank test addresses paired comparisons between algorithms, while the Friedman test manages problem-specific grouping effects. For multiple algorithm comparisons, we used Holm’s alpha correction to control error rates (IsmailFawaz2018deep, ; brockhoff2023gecco, ; picard2024untrained, ).
Algorithm Runtime Recording
Each algorithm is timed for the total time it takes to run one experiment trial. We take the overall mean over the 20 trials and over all benchmark problems, resulting in a single value average time ($t_{\text{avg}}$) to represent the average runtime of each algorithm.
5 Results and Evaluations
In this section, we report the algorithm performance on a subset of the 63 benchmark optimization experiments and the ablation study on GIT-BO and vanilla TabPFN v2 BO. The complete optimization results for all synthetic and real-world engineering problems are detailed in the appendix.
Overall Statistical Ranking and Algorithm Runtime Tradeoffs
Across all problem variants, Figure 2 (a) shows that GIT-BO achieves the top statistical performance rank based on the optimization results (1.97), followed by SAASBO (3.51), ALEBO (d=10) (4.11), and HESBO (d=20) (5.14). Even when we enforce a strict iteration limit at 100 iterations, GIT-BO still outperforms other GP-based algorithms in the statistical rank demonstrated in Figure 2 (b). A detailed rank-iteration evolution curve in Figure 2 (c) shows that GIT-BO becomes dominant within the first 25 iterations and keeps that lead. Figure 2 (d) plots the average run time $t_{\text{avg}}$ of each algorithm versus the statistical rank. GIT-BO requires $\approx 10^{3}$ sec per trial, two orders of magnitude faster than SAASBO and ALEBO. With the first place in rank and third place in $t_{\text{avg}}$, GIT-BO is the Pareto frontier of speed and quality.
Figure 2: (a) and (b): Statistical ranking of the overall performance across 63 benchmark problems. GIT-BO ranked the top at optimization results at both fixed 100- and 200-iteration evaluations. (c): Plot of average algorithm rank at each iteration. Due to the compute limit, we only run ALEBO and SAASBO for 100 iterations. We compare their final results after 100 iterations with other algorithms run for 200 iterations. The plot shows that GIT-BO achieves a fast convergence at ranking performance as it remains the top of all algorithms after 25 iterations. (d): Plot of average time vs overall statistical rank. The shorter time and smaller rank perform better (bottom left corner), so we show GIT-BO at the Paerto front as the best algorithm.
Scalable Synthetic Benchmarks (100 $\leq$ D $\leq$ 500)
Overall, GIT-BO outperforms other SOTA algorithms in 35 out of 45 variants of the scalable synthetic benchmark problems. The convergence results for these scalable synthetic problems can be summarized in three categories in Figure 3. First, GIT-BO has absolute domination over several problems (e.g., Ackley), outperforming all other algorithms in all dimensions. The second category, the most common observed ones, are that ALEBO (d=10) or SAASBO match or outperform GIT-BO in $D=100$ ; however, GIT-BO maintains its top optimization efficiency in the higher dimensions $D\geq 300$ (e.g., Levy). Finally, GIT-BO struggles with very few problems (e.g., Michalewicz), where the current SOTA methods with the trust region and linear embedding finding strategies dominate.
Figure 3: Optimization results on the synthetic benchmarks comparing our method against baseline algorithms. The solid line represents the median best function value achieved over 20 trials, with shaded regions indicating the 95% confidence interval. GIT-BO ranks the top for nine variants out of 15, showing better performance at higher dimensional problems and faster convergence for the first 100 iterations. GIT-BO also has a lower variance compared to other baselines. In practice this means GIT-BO not only converges faster in expectation but also delivers reliably strong results without requiring multiple restarts, reducing both computational cost and risk. Full statistical tests and per-problem plots are provided in the Appendix.
Real-world Problem
Figure 4 illustrates the convergence results for a subset of eight real-world problems. GIT-BO outperforms all baselines on the four 100+D CEC engineering tasks on power system optimization and the 388D SVM hyperparameter optimization problem, showing its great potential to tackle real-world applications. GIT-BO achieves the second and third best in the LassoDNA problem and car problems (MOPTA08 Car and Mazda), respectively. One possible explanation is that the distribution of these engineering benchmarks might be very different from the training distribution of our TFM, TabPFN v2. Although SAASBO achieves the best for these problems, it takes on average three hours to solve them, while the second best, GIT-BO, takes only fifteen minutes. This showcases that GIT-BO offers a compelling trade-off for resource-constrained engineering optimization tasks where time-to-solution is critical.
Figure 4: Optimization results on the engineering benchmarks comparing our method against baseline algorithms. The solid line represents the median best function value achieved over 20 trials, with shaded regions indicating the 95% confidence interval. GIT-BO performs the best for the CEC problems, and maintains a 1.15 average ranking over the real-world problem set.
Ablation Study
To rigorously assess the necessity of gradient-informed subspace exploration, we conducted an ablation study comparing GIT-BO with a variant employing the vanilla TabPFN v2 model without adaptive gradient-informed subspace identification. Two acquisition functions are tested with the vanilla TabPFN v2: expected improvement acquisition (EI) (used in the original PFNs4BO (muller2023pfns, )) and Thompson Sampling (TS), which is used by GIT-BO. The convergence plot results shown in Figure 5 clearly indicate that both vanilla TabPFN v2 with EI and TS has failed the high-dimensional search significantly compared to GIT-BO. Specifically, we observe a drastic degradation in convergence speed and final optimization outcomes. This confirms our hypothesis that vanilla TabPFN v2, regardless of the selection of the acquisition function, fails to capture crucial directions in high-dimensional spaces without the assistance of the GI subspace. This reinforces the critical role of our proposed gradient-based refinement strategy in achieving robust performance.
Figure 5: Ablation studies of vanilla TabPFNv2 BO with GIT-BO. GIT-BO consistently outperform other algorithms without GI subspace search across all tested problems (8.6 times better regrets), indicating GIT-BO is the best approach and vanilla TabPFNv2 is not an ideal candidate for performing high-dimensional BO.
6 Discussion
Novelties and Strengths
Our results demonstrate that a tabular foundation model can serve as a top-class surrogate in high-dimensional Bayesian optimization. By combining TabPFN’s one-shot regression with a data-driven, gradient-informed subspace that is updated after every evaluation, GIT-BO converges faster than carefully tuned Gaussian-process baselines on sixty-three benchmarks, maintaining its lead even as dimensionality grows to 500D. The approach removes the computational bottleneck of GP retraining and latent space searching while preserving the exploration–exploitation balance that characterizes principled BO (two orders of magnitude higher than SAASBO). Taken together, these properties position GIT-BO as an accessible “drop-in” optimizer for engineers who already rely on large language models and are now seeking equally powerful tools for structured data and design optimization.
Limitations
Our approach, although effective, faces several practical limitations. First, the computational requirements of TabPFN present significant hardware constraints. As a large foundation model, TabPFN requires GPU acceleration for efficient inference. This limits accessibility for users without GPUs. Second, the current implementation imposes a hard dimensionality cap of 500D, as the foundation model was trained specifically with this input dimension limit. Problems exceeding this dimensionality cannot be directly addressed without architectural modifications or using ensembles. Expanding to higher-dimensional spaces would require retraining the entire foundation model with an expanded feature capacity, representing a substantial computational investment. Third, although GIT-BO is largely parameter-free, the rank $r$ of the GI subspace and the choice of reference point $x_{\text{ref}}$ for back-projection could influence the algorithm’s performance. We provide an initial sensitivity analysis in the Appendix. However, future work will explore using an automated heuristic or meta-learning of these hyperparameters.
Societal Impact
GIT-BO advances high-dimensional BO by leveraging gradient-informed subspace identification from pre-trained tabular foundation models, significantly improving optimization robustness and scalability. This methodological innovation can empower researchers and practitioners to efficiently tackle previously challenging optimization problems in critical application areas of engineering design where objective evaluation is expensive (wang2023high, ; deneault2021toward, ; lam2018advances, ). Additionally, the pre-trained nature of foundation models enables effective reuse across diverse tasks, potentially lowering computational overhead by eliminating training application-specific prediction models in repeated optimization studies. This especially benefits adaptive experimentation, such as drug and material discovery, where the research results and the researcher’s schedule depend on the optimization algorithm’s output (wang2022bayesian, ; griffiths2020constrained, ; shields2021bayesian, ). Thus, GIT-BO not only democratizes high-dimensional Bayesian optimization methodologies but also contributes indirectly to reducing resource consumption in long-term scientific research and industrial optimization cycles.
Conclusion and Future Work
In this paper, we introduced Gradient-Informed Bayesian Optimization using Tabular Foundation Models (GIT-BO), a method integrating the computational efficiency of tabular foundation models with an adaptive, gradient-informed subspace identification strategy for Bayesian optimization. Our comprehensive evaluations across synthetic and real-world benchmarks demonstrate that GIT-BO consistently surpasses state-of-the-art Gaussian process-based methods in high-dimensional optimization scenarios (up to 500 dimensions). By eliminating iterative surrogate retraining and dynamically identifying critical subspaces, GIT-BO also accelerates convergence and reduces computational costs by orders of magnitude compared to traditional GP methods.
The capability to efficiently and effectively handle previously intractable high-dimensional optimization tasks broadens the practical applicability of Bayesian optimization. This advancement is particularly impactful for automated machine learning pipelines, scientific experimentation, and complex engineering design problems where dimensionality has historically impeded efficient optimization. As tabular foundation models continue to grow in model size and improve the prediction accuracy, we anticipate even greater optimization performance from approaches like GIT-BO, further extending Bayesian optimization to even more challenging real-world scenarios. Future work can leverage our method in multiple directions, such as constrained optimization, mixed-integer problems, and multi-objective optimization, through modifications to the acquisition functions and embedding techniques.
Appendix A Additional Details on Benchmark Algorithms and Problems Implementations
A.1 Reproducibility Statement
The code for experiments and the GIT-BO algorithm is provided in the supplemental material zip file. It can also be found in this anonymous link: https://anonymous.4open.science/r/GITBO-9EFE/README.md.
A.2 GIT-BO Implementation Details
The GIT-BO algorithm was implemented using Python 3.12 with the TabPFN v2.0.6 implementation and model (link: https://github.com/PriorLabs/TabPFN and https://huggingface.co/Prior-Labs/TabPFN-v2-reg, license: Prior Lab License (a derivative of the Apache 2.0 license (http://www.apache.org/licenses/))). Instead of using the off-the-shelf TabPFN code for GIT-BO, we make a wrapper that converts the numpy calculations to PyTorch for torch.backward() gradient calculations. The system uses BoTorch v0.12.0 and PyTorch 2.6.0+cu126 for the underlying optimization framework. For the gradient-informed subspace identification, a rank of 15 was selected for the GI subspace matrix. All experiments were conducted on a GNU/Linux 6.5.0-15-generic x86_64 system running Ubuntu 22.04.3 LTS as the operating system, ensuring a consistent computational environment across all benchmark tests.
A.3 GP-based Benchmark Algorithms Implementation Details
We benchmark GIT-BO against four high-dimensional BO methods, including SAASBO (eriksson2021high, ), TURBO (eriksson2019global, ), HESBO (nayebi2019framework, ), and ALEBO (letham2020re, ).
As most BO code is archived in the BoTorch or Ax Platform [^1] https://github.com/facebook/ax library, we use Botorch 0.12.0 for SAASBO and TURBO. However, since Ax Platform has officially deprecated ALEBO, we have created another environment to run ALEBO using the previous version of Ax Platform and related packages that are compatible with the original code of ALEBO.  [^2] https://github.com/facebook/Ax/issues/1924 The environment setups are detailed in the provided code zip file.
A.4 Benchmark Problems Implementaion Details
The source and license details of our benchmark problems are described in the following paragraphs.
Synthetic Problems:
The implementations for the nine synthetic functions are taken from Botorch (balandat2020botorch, ) (link: https://github.com/pytorch/botorch/blob/main/botorch/test_functions/synthetic.py, license: MIT license, last accessed: May 1st, 2025). The bounds of each problem are the default implementation in Botorch. Detailed equations for each problem can be found here: https://www.sfu.ca/~ssurjano/optimization.html.
CEC2020 Power System Problems:
We examine a subset of six problems, specifically those with design spaces exceeding 100 dimensions, from the CEC2020 test suite (kumar2020test, ) (link: https://github.com/P-N-Suganthan/2020-Bound-Constrained-Opt-Benchmark, license: no licence, last accessed: May 1st, 2025). The code is initially in MATLAB, and we translate it into Python, running pytest to ensure the implementations are correct. While these problems incorporate equality constraints ($h_{j}(x)$), they are transformed into inequality constraints ($g_{j}(x)$) using the methodology outlined in the original paper (kumar2020test, ), as constraint handling is not the primary focus of this research. These transformed constraints are subsequently incorporated into the objective function $f(x)$ as penalty terms.
$g_{j}(x)=|h_{j}(x)|-\epsilon\leq 0\;,\;\epsilon=10^{-4}\;,\;j=1\sim C$
$f_{penalty}(x)=f(x)+\rho\sum^{C}_{j=1}max(0,g_{j}(x))$
We set a different $\rho$ penalty factor for each problem, respectively, to make the objective and constraint values have a similar effect on $f_{penalty}(x)$.
Table 1: CEC2020 Benchmark Problems Penalty Transform Factor $\rho$
Problems CEC34 CEC35 CEC36 CEC37 CEC38 CEC39
$\rho$ 0.01 0.0002 0.001 0.04 0.02 0.04
LassoBench Benchmarks:
The implementation is taken from the original implementation (vsehic2022lassobench, ) (link: https://github.com/ksehic/LassoBench, license: MIT and BSD-3-Clause license, last accessed: May 1st, 2025).
Rover:
The implementation is taken from (wang2018batched, ) (link: https://github.com/zi-w/Ensemble-Bayesian-Optimization, license: MIT license, last accessed: May 1st, 2025).
MOPTA08 Car:
The MOPTA08 executables are taken from the paper (papenmeier2022increasing, )’s personal website (link: https://leonard.papenmeier.io/2023/02/09/mopta08-executables.html, license: no license, last accessed: May 1st, 2025). The MOPTA08 Car’s penalty transformation follows the formation of (eriksson2021high, )’s supplementary material.
Mazda Bechmark Problems:
The implementation is taken from (kohira2018proposal, ) (link: https://ladse.eng.isas.jaxa.jp/benchmark/, license: no license, last accessed: May 1st, 2025). The Mazda problem has two raw forms: a 4-objectives problem 148D (Mazda_SCA) and a 5-objectives  [^2] D problem (Mazda), and both of them have inequality constraints. For both problems, we equally weight each objective to form a single objective and perform a penalty transform:
$f_{multiobj\_penalty}(x)=\frac{1}{N}\sum^{N}_{i=1}f(x)+\rho\sum^{C}_{j=1}max(0,g_{j}(x))$
where $N$ is the number of objectives, $C$ is the number of inequality constraints, and we use $\rho=$ 10 for both variants of Mazda problem.
Table 2 summarizes the type of problems and their respective tested dimensions.
Table 2: High-Dimensional Benchmark Problems
Problems Source Type Dimension ($D$) Tested
Ackley Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Dixon-Price Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Griewank Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Levy Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Michalewicz Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Powell Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Rastrigin Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Rosenbrock Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
Styblinski-Tang Botorch (balandat2020botorch, ) Synthetic 100, 200, 300, 400, 500
CEC34 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 118
CEC35 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 153
CEC36 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 158
CEC37 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 126
CEC38 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 126
CEC39 CEC2020 Benchmark Suite (kumar2020test, ) Real-World 126
LassoSyntMedium LassoBench (vsehic2022lassobench, ) Synthetic 100
LassoSyntHigh LassoBench (vsehic2022lassobench, ) Synthetic 300
LassoDNA LassoBench (vsehic2022lassobench, ) Real-World 180
Rover Previous BO studies (wang2018batched, ; xu2024standard, ) Real-World 100, 200, 300, 400, 500
MOPTA08 CAR Previous BO studies (eriksson2019global, ; papenmeier2022increasing, ) Real-World 124
MAZDA Mazda Car Bechmark (kohira2018proposal, ) Real-World 222
MAZDA SCA Mazda Car Bechmark (kohira2018proposal, ) Real-World 148
SVM Previous BO studies (eriksson2021high, ; xu2024standard, ; papenmeier2022increasing, ) Real-World 388
Appendix B Parameter Sweep
We conducted a parameter sweep to evaluate the sensitivity of GIT-BO’s optimization performance to the chosen dimensionality ($r$) of the gradient-informed active subspace. The results are presented in Figure 6. The empirical findings suggest that very high-dimensional subspaces (e.g., $r=40$ for 100-dimensional problems) result in significant performance degradation due to diminished effectiveness of the gradient-informed search direction. Conversely, lower-dimensional subspaces generally showed superior performance. Although problem-specific hyperparameter optimization would likely enhance these results, we maintained fixed parameter values across ($r=15$) all benchmarks to ensure fair comparison, demonstrating that GIT-BO consistently achieves top-ranked performance on average.
Figure 6: Performance comparison across various gradient-informed subspace dimensions ($r$) in GIT-BO. Median optimization results across 20 trials are shown, with shaded regions depicting 95% confidence intervals. High-dimensional subspaces (e.g., $r=40$) exhibit notable performance deterioration, and lower subspace dimensions tend to perform consistently better.
Figure 7: Rank of the final optimization solution performance comparison across various gradient-informed subspace dimensions ($r$) in GIT-BO.
Appendix C Full Results
We present the comprehensive optimization results for all benchmark problems investigated in the main paper. Figures 8 to 11 provide the complete convergence curves for synthetic and real-world benchmarks, respectively. The results consistently demonstrate GIT-BO’s robust performance across different classes of high-dimensional problems, clearly indicating its advantage in both convergence speed and final solution quality compared to baseline methods. Notably, in Figure 12, the statistical ranking of the variance in the final results for 20 trials shows that GIT-BO has a lower variance than other baselines, which means that GIT-BO converges faster in expectation and delivers reliably strong results without requiring multiple restarts, practically reducing computational cost and risk.
Figure 8: Full convergence curves on engineering benchmarks illustrating the median best solutions obtained across multiple trials. Confidence intervals (95%) are shown as shaded areas around the median lines. The minimal median best solution among all algorithms at the final iteration is marked in “x”. The results demonstrate that SAASBO excels in the Rover and car (MOPTA08 and Mazdas) problems, and GIT-BO leads the CEC power system optimization problem. Though GIT-BO is second-best for real-world problems, engineers might benefit from its 100x faster runtime if time is a hard constraint.
Figure 9: Violin plot statistical summary of final optimization performance for engineering functions, summarizing results over 20 trials.
Figure 10: Detailed optimization results for synthetic benchmarks demonstrating the relative performance of GIT-BO versus competing algorithms. Median best-found function values are illustrated by solid lines, with shaded bands denoting 95% confidence intervals. The minimal median best solution among all algorithms at the final iteration is marked in “x”. Results highlight the robustness of GIT-BO in achieving superior performance across diverse synthetic problems, especially in challenging higher dimensional settings ($D>$300). The only two exceptions that GIT-BO fails dramatically in all $D$s are Styblinski-Tang and Michalewicz, which can potentially be explained by the “No Free Lunch” theorem (wolpert1997no, ).
Figure 11: Violin plot representations of final optimization performance for synthetic functions, summarizing results over 20 trials.
Figure 12: Statistical ranking of the overall variance across (a) Real-World and (b) Synthetic benchmark problems. GIT-BO consistently ranks at the top, showing the lowest variance during optimization convergence.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the abstract gives a brief summary of the scope, and the introduction section lists out the contributions. See Abstract and Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 6 paragraph “Limitations".
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 3 and Algorithm 1 provide the overview and the pseudo code for the GIT-BO algorithm, and the Appendix contains the implementation details of GIT-BO and the benchmark algorithms. The source code for using the algorithms and running the experiments is provided in the supplemental material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All source code, environment files, and run scripts needed to reproduce every experiment are packaged in an anonymized zip in the supplemental material. We will migrate this archive to a public GitHub repository under an open-source license upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 3 and Section 4 details the GIT-BO’s hyperparameter selection and experiment testing processes. There is no training information since the algorithm does not involve any model training.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: As stated in Section 4, we repeat optimization runs twenty times for each optimization algorithm and perform statistical ranking and evaluations of the optimization results for fair comparison. Section 5 and the Appendix include the results of the statistical ranking of the optimization algorithm convergence plots with the 95% confidence interval shown.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See section 4 “Algorithm tests”.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, this paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Section 6 paragraph “Societal Impacts”.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: See Section 6 paragraph “Societal Impacts”.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Appendix A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All new algorithm code introduced in the paper and its documentation are included in an anonymized zip file in the supplementary material. They will be published on a public GitHub under an open-source license upon acceptance.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 GIT-BO consistently ranks at the top, showing the lowest variance during optimization convergence.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Yes, the abstract gives a brief summary of the scope, and the introduction section lists out the contributions. See Abstract and Section 1.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See Section 6 paragraph “Limitations".
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Section 3 and Algorithm 1 provide the overview and the pseudo code for the GIT-BO algorithm, and the Appendix contains the implementation details of GIT-BO and the benchmark algorithms.
-The source code for using the algorithms and running the experiments is provided in the supplemental material.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: All source code, environment files, and run scripts needed to reproduce every experiment are packaged in an anonymized zip in the supplemental material.
-We will migrate this archive to a public GitHub repository under an open-source license upon acceptance.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: Section 3 and Section 4 details the GIT-BO’s hyperparameter selection and experiment testing processes.
-There is no training information since the algorithm does not involve any model training.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: As stated in Section 4, we repeat optimization runs twenty times for each optimization algorithm and perform statistical ranking and evaluations of the optimization results for fair comparison.
-Section 5 and the Appendix include the results of the statistical ranking of the optimization algorithm convergence plots with the 95% confidence interval shown.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: See section 4 “Algorithm tests”.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Yes, this paper conforms, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: See Section 6 paragraph “Societal Impacts”.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: See Section 6 paragraph “Societal Impacts”.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: See Appendix A.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64.
-Justification: All new algorithm code introduced in the paper and its documentation are included in an anonymized zip file in the supplementary material.
-They will be published on a public GitHub under an open-source license upon acceptance.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.